OpenStudy (anonymous):

A game of tug of war takes place, and the team on the right side pulls with 7.50 x 10^3 N of force, while the team on the left pulls with 6.90 x 10^3 N. The team on the right moves the rope a distance of 2.00 m. There are 8 team members on each team. How much work did each individual team member do?

OpenStudy (mayankdevnani):

do you know that the work on each side is equal?

OpenStudy (mayankdevnani):

@theequestrian

OpenStudy (anonymous):

Yes

OpenStudy (mayankdevnani):

so , \[\huge work~done=force \times displacement\] where, \[\huge F=7.50 \times 10^3 N\] and \[\huge S=2m\]

OpenStudy (mayankdevnani):

so what is the work done by right team???

OpenStudy (anonymous):

1.5 * 10^4

OpenStudy (anonymous):

In Joules

OpenStudy (anonymous):

Oh & the left side is equal to the right! So my answer is 1.5*10^4. Wow I feel really silly right now.. But thank you!

OpenStudy (mayankdevnani):

welcome:)

OpenStudy (mayankdevnani):

and then divide this by 8 and you get your answer

OpenStudy (anonymous):

Oh right!! :)

OpenStudy (mayankdevnani):

because you need to find the work by each.

OpenStudy (mayankdevnani):

i think you can solve it. @theequestrian

OpenStudy (anonymous):

It would be 1875J

OpenStudy (mayankdevnani):

hey!!! it want to calculate...but i trust on you..... i think you are right

OpenStudy (anonymous):

Haha, thank you:)

OpenStudy (mayankdevnani):

welcome:)